OpenStudy (daniellelovee):

A student reduces the temperature of a 300 cm3 balloon from 60°C to 30°C. What will the new volume of the balloon be? (Remember to be careful with the units.)

OpenStudy (daniellelovee):

300 cm3 273 cm3 150 cm3 600 cm3

OpenStudy (daniellelovee):

I got 600 is this correct? @Somy (sorry girl or boy I know I'm annoying xD )

OpenStudy (daniellelovee):

@Somy

Parth (parthkohli):

(Remember to be careful with the units.)

OpenStudy (alekos):

volume is proportional to temperature so if the temp goes down then so does the volume

OpenStudy (somy):

V1/T1 = V2/T2 V unit is m3 T unit is Kelvin

OpenStudy (somy):

use the formula and again make sure your units are correct

OpenStudy (somy):

V1 = 300 cm3 which needs to be changed to m3 T1 = 60 °C which needs to be converted to Kelvin (you can convert it by adding 273 to Celsius value) T2 = 30 °C needs to be converted to Kelvin also V2 = X m3 so use the formula your answer will be in m3 so make sure to convert it to cm3 to get the answer to match unit of options given (p.s. dont worry, i'll help if i can :3 doesnt matter how many times, so its not annoying or anything)

OpenStudy (alekos):

Good work Somy but I think V2 will be in cm3

OpenStudy (somy):

@alekos the formula is better to be used with its appropriate units which are Kelvin for temperature and m3 for volume so she/he has to convert her/his cm3 value of given V1 to m3 to avoid any error and final answer for V2 will be in m3 which then she/he can convert to cm3 and be sure there is no unit error in his/her calculation

OpenStudy (alekos):

OK fair enough. MKSA units are always better.

Parth (parthkohli):

@Somy It doesn't actually make any difference if you know what you're doing. \(\rm 1 m^3 = 10^6 cm^3\), so you can just keep working with cubic centimeters because a cubic centimeter can directly be converted to a cubic meter.

OpenStudy (somy):

i know but thats me and you who can already do this since we have done a lot of it but i dont know how many of these type of questions has @Daniellelovee done which is why to be on the safe side i think its better for him/her to use SI unit system for this very equation just to avoid error

Parth (parthkohli):

i getcha. it's generally good practice to be careful.

OpenStudy (dan815):